If 25.6 g of NO and 26.9 g of O₂ react together, how many grams of NO₂ can be formed via the reaction below? 2 NO (g) + O₂ (g) → 2 NO₂ (g)
Answer #1


Molar mass of NO,
MM = 1*MM(N) + 1*MM(O)
= 1*14.01 + 1*16.0
= 30.01 g/mol


mass(NO)= 25.6 g

use:
number of mol of NO,
n = mass of NO/molar mass of NO
=(25.6 g)/(30.01 g/mol)
= 0.853 mol

Molar mass of O2 = 32 g/mol


mass(O2)= 26.9 g

use:
number of mol of O2,
n = mass of O2/molar mass of O2
=(26.9 g)/(32 g/mol)
= 0.8406 mol
Balanced chemical equation is:
2 NO + O2 ---> 2 NO2


2 mol of NO reacts with 1 mol of O2
for 0.853 mol of NO, 0.4265 mol of O2 is required
But we have 0.8406 mol of O2

so, NO is limiting reagent
we will use NO in further calculation


Molar mass of NO2,
MM = 1*MM(N) + 2*MM(O)
= 1*14.01 + 2*16.0
= 46.01 g/mol

According to balanced equation
mol of NO2 formed = (2/2)* moles of NO
= (2/2)*0.853
= 0.853 mol


use:
mass of NO2 = number of mol * molar mass
= 0.853*46.01
= 39.25 g
Answer: 39.2 g